OpenStudy (anonymous):

The perimeter of a rectangle is 200 feet. Describe the possible lengths of a side if the area of the rectangle is not to exceed 900 feet. ANSWER: The length (in feet) of a side is in the interval ___ or in the interval ____

OpenStudy (anonymous):

Need my interpretation??

OpenStudy (anonymous):

yes!!! very bad

OpenStudy (anonymous):



OpenStudy (anonymous):

area = W*L < =900

OpenStudy (anonymous):

from the area, you have 2 numbers, W and L , maximum product of them is 900. Now consider what happens 900 = 1*900 --> the W =1 and L=900. However 1+900 =901 , too much bigger than W+L =100 above. Got me so far??

OpenStudy (anonymous):

900 = 2*450 also. The same reason of above, we reject this result because 2+450=452 \(\neg\)100

OpenStudy (anonymous):

how bout 900=10*90 ?? which gives us the W =10 and the L = 90 Okie dokie, If W =10, L =90, then W+L =10+90= 100 and W*L=900 so, the first solution is 10 for width and 90 for length. got me so far??

OpenStudy (anonymous):

yes yes yes

OpenStudy (anonymous):

ok, the leftover is your duty. hihihi I don't have any words to interpret more. :)

OpenStudy (anonymous):

wait so if those are your solutions than how do you write those in interval notation? That's the part that also stumps me

OpenStudy (anonymous):

Actually, if the condition of perimeter is lesser or equal 100, then you have interval. In this case, you have just = 100, you don't have interval, just exactly 1 solution. Hopefully you missed the "lesser" part on it.

OpenStudy (anonymous):

*200, not100

OpenStudy (anonymous):

the question is asking me for two intervals

OpenStudy (anonymous):

with two blanks

OpenStudy (anonymous):

The length in feet of a side is the interval ___ or in the interval ____

OpenStudy (anonymous):

I know!! but the condition of the perimeter must be "lesser or equal 200" not just "is 200

OpenStudy (anonymous):

right right

OpenStudy (anonymous):

so would that make one of them (-inf,200]

OpenStudy (anonymous):

for example, if I have W =45, L = 20, I still have 45*20=900 but 45+20 =65 not 100

OpenStudy (anonymous):

help me please @tkhunny

OpenStudy (tkhunny):

Just thinking about the Area \(0 < Length < 900\) \(0 < Width < 900\) Now to consider the perimeter. \(0 < Length < 100\) \(0 < Width < 100\) Now the relationship between the two. L*W < 900 2L+2W = 200 ==> 2L + 1800/L = 200 ==> 2L^2 - 200L + 1800 = 0 L = 90 or 10 <== Are these the limits? Try it out. L = 85, then W = 900/85 = 4.615 and 2*85 + 2*10.588 = 191.176 -- Hmmm... What are we missing? gtg

OpenStudy (anonymous):

That's more professional than my interpreting. However, the condition of the perimeter should be changed to < = 200. :)

OpenStudy (anonymous):

Okay I understand how to solve it now I just don't know how to write the answer down!

OpenStudy (anonymous):

Thank you for that step by step though from both of you that helped a lot

OpenStudy (tkhunny):

Yup. Go OOOPS! It hit me about an hour later that I was using Area = 900 and Perimeter < 200 instead of the other way around.

OpenStudy (anonymous):

Can somebody please tell me how to write the answer in interval notation please